# Chained computerized adaptive testing for the Force Concept Inventory

### v1

@@ Abstract
Although conceptual assessment tests are frequently administered in a pre/post-semester fashion, there are inherent issues with this paradigm. Specifically, education researchers and instructors have limited ability to observe the progression of student conceptual understanding throughout the course. Furthermore, instructors are limited in the usefulness of the feedback they can give to the students involved. To address these issues, we propose the use of computerized adaptive testing (CAT) and increasing the frequency of CAT-based assessments during the course, while reducing the test length per administration, thus keeping or decreasing the total number of test items administered throughout the course. The feasibility of this idea depends on how far the test length per administration can be reduced without compromising the test accuracy and precision. Specifically, the overall test length is desired to be shorter than when the full assessment is administered as a pretest and subsequent post-test. To achieve this goal, we developed a CAT algorithm that we call Chain-CAT. This algorithm sequentially links the results of each CAT administration using collateral information. We developed the Chain-CAT algorithm using the items of the Force Concept Inventory (FCI) and analyzed the efficiency by numerical simulations. We found that collateral information significantly improved the test efficiency, and the overall test length could be shorter than the pre-post method. Without constraints for item balancing and exposure control, simulation results indicated that the efficiency of Chain-CAT is comparable to that of the pre-post method even if the length of each CAT administration is only 5 items and the CAT is administered 9 times throughout the semester. However, when these constraints are imposed, we found that the efficiency of the Chain-CAT with the test length of 5 items is lower than that of the pre-post method. Based on this result and analysis of which items were administered, we suggest expanding the FCI item bank by creating additional items or incorporating items from other research-based assessments with high discrimination parameters.
@@ I Introduction
The pre-post method is a standard method to analyze the effectiveness of instruction, but there are limitations to the approach. First, the pre- and post-test results provide only snapshots of students’ understanding at the beginning and end of the course, limiting the ability to observe the progression of conceptual understanding throughout the duration of the course. Second, students may see little benefit in their taking the assessment when the focus of the assessment is to reflect on the year’s instruction and improve instructional practices for future students, with no feedback to the students who actually take the assessment. This situation can make students feel burdened and may decrease their engagement and motivation to take the assessment.
To address these issues associated with the pre-post paradigm, we propose to increase the frequency of the assessments during the course while decreasing the test length per administration, thus maintaining or decreasing the total number of test items during the course. This idea is inspired by the microgenetic method [11, 12, 13, 14, 15, 16, 17, 18], which involves frequent quantitative and/or qualitative measurements. The microgenetic method involves 1) sampling data at a frequency that is assumed to be high compared to the rate of change of the phenomenon of interest and 2) collecting data for the entire period of change [17], allowing teachers to monitor the progression of students’ understanding. Although these methods are generally implemented with small student cohorts, Sayre and Heckler [13, 14] used the microgenetic approach in frequently administering simple multiple choice conceptual questions to a large number of students.
Our idea of frequent administration of short assessments can be used as a form of formative assessment. Providing automated feedback according to each student’s set of responses is expected to increase the usefulness of the survey for students and reduce their sense of burden.
The feasibility of the above idea depends on how far the test length per administration can be reduced. A reference value of the total test length during a course is 60 items, as this is the total test length if the full FCI is administered both as the pre- and post-test. If the course involves 10 administrations of the shortened FCI, it is then desired that the test length per administration would be less than 6 items.
@@
We utilize a CAT algorithm that takes advantage of collateral information [31, 19]. Collateral information is the relevant empirical information on the respondents, for example, age, grade, or previous test scores. This information can be used to select the first item in CAT, accelerating the convergence of the proficiency estimation, and hence improving test efficiency [19]. Specifically, as we describe below, we use the proficiency estimate of each respondent in a test for selecting items and estimating respondent proficiency level in the next test for the respondent. Since this CAT algorithm sequentially links the test result of each administration, we call this algorithm Chain-CAT.
@@
Several researchers have studied the development of a CAT-based formative assessment system. Veldkamp et al. [32] discussed the viability of a computerized adaptive learning system which responds to students based upon their test results. Choi and McClenen [33] developed and analyzed the validity and efficiency of an adaptive formative assessment system based on CAT with the time series proficiency estimation utilizing Bayesian networks. Yang et al. [34] developed a formative assessment system based on CAT and learning memory cycle. While there have been previous studies such as these, our goal is to construct an algorithm from our novel design of chaining collateral information as described above and to put it to practical use in the context of physics education research.
The objective of this study is to examine the feasibility of the Chain-CAT version of the FCI (Chain-CAT FCI). Specifically, our research questions are: (i) How much does collateral information improve the test efficiency when it is used to sequentially link the results of each CAT administration? (ii) Is it possible to shorten the total test length of Chain-CAT FCI in a course to be less than that of the full FCI conducted as the pre- and post-test (60 items)? Under what conditions is this possible? (iii) Do the 30 FCI items comprise a sufficient item bank for the effective use of Chain-CAT? If not, what improvement is necessary? The focus of this paper is on the analysis of the feasibility of the Chain-CAT FCI using numerical simulations. The development of feedback methods to deliver formative assessment will be the subject of future work.
@@ II.1 Item response model and item bank
We estimated the item parameters using the FCI survey data obtained in our previous study [40, 41]. In that study, we administered the full-length paper-and-pencil FCI in class to 2882 Japanese university students from April 2015 to April 2018. From this dataset, we removed aberrant responses to be left with 2712 valid responses. We confirmed that the standard errors of the parameter estimates are not significant. The assumptions of unidimensionality, overall local independence, and goodness of fit were confirmed in this study [40, 41], which validates the use of the 2PL model in our analysis. Specifically, we examined the unidimensionality of the FCI via a principal component analysis with the tetra-choric correlation matrix [39], we evaluated the local independence assumption by using Yen’s statistic [42], and we evaluated the goodness of fit of the 2PL model to the response data with the standardized root mean square residual [43].
@@ II.4 Improvement of the validity of Chain-CAT
Second, we controlled the item exposure for the Chain-CAT FCI. Generally, if the items in an item bank are very informative at the average ability level, they are selected more frequently than the other items [20]. Although a given item does not appear more than once in a given test, the simple Chain-CAT algorithm can result in that item frequently appearing across the tests. If students repeatedly answer certain items each week, they may memorize the content of the items they solved the previous week and answer them without rethinking, which may not reflect their actual proficiency. Therefore, we implemented an algorithm that limits the number of times an item can be administered in a course. Specifically, we used an algorithm whereby once an item is administered a specific number of times in a course, it is removed from the list of items that can be administered (described in more detail below).
@@ II.5 Numerical simulation procedure
We conducted a numerical simulation to analyze the efficiency of the Chain-CAT and to search for the optimal settings. The outline of the procedure is as follows. (1) Assume a progression model of the true-value proficiency as a function of time (the unit is taken to be 1-2 weeks) in a course. (2) Generate response data for a given true value of based on the item response model and calculate the estimates of . (3) Calculate the accuracy and precision by the root-mean-square error (RMSE) from the true and the estimated values of and compare it to a reference value. In what follows, we explain the details of the above procedure.
@@ II.5.1 Progression models of the true proficiency level
To analyze the efficiency of the Chain-CAT as generally as possible, we assumed various progression patterns of the true-value proficiency . Specifically, we assumed the following progression models for : a) a stationary model in which is constant in time, b) a linear model in which increases linearly over time, and c) a step model in which increases significantly only at a certain point in time. Although it is uncommon in introductory physics courses for student scores on the FCI to decrease, we confirmed that the simulation results of the case when decreases linearly or decreases significantly only at a certain point are similar to the results of b) or c), respectively. We supposed that most of the student’s progressions are one of these patterns or a combination of them. Moreover, we analyzed the test efficiency of the Chain-CAT in various proficiency ranges for each model. For example, in the stationary model, we analyzed the efficiency in the middle students and the high students . We confirmed that the low range yields similar results to the high range , so we do not show the result below. We fixed the number of tests conducted throughout the simulated course to be 9 times. This would be a reasonable number of tests to be administered in a course consisting of 10-15 weeks.
@@ II.5.2 Generation of the response data
We generated response data based upon a Monte Carlo simulation, which is commonly used in the development of CAT [46]. Specifically, we generated the responses for Chain-CAT using the function simulateRespondents of the catR package. In the simulation, the response to a selected item is generated based on the calibrated item parameters of that item and the value of true proficiency at a point in time. For example, in Eq. (1), suppose that the probability of a correct response for an item is calculated to be 0.75 for a given true proficiency. A random number is generated from a uniform distribution within a range of 0 to 1. If the value is 0.75 or less, the generated response is coded as “correct” for that item. If the value is greater than 0.75, then the generated response is “incorrect.” The EAP method is used to estimate the proficiency of the respondent, an item selection method (e.g., MFI criterion) is then used to choose the next item based upon that estimated proficiency, and the process repeats until reaching the predetermined test length. Specifically, we varied the test length to be 5, 10, and 15. Then, for each length, we calculated accuracy and precision. To summarize, provided calibrated item parameters and a given true proficiency level, the simulateRespondents function generates an entire set of correct/incorrect responses and calculates estimated proficiencies for a given test length of each Chain-CAT administration. This is done for each time point in a course. In this analysis, we generated 10 000 response data based on the same true theta at a given moment in time.
@@ II.5.3 Algorithm to improve the validity of Chain-CAT
As we mentioned above, to improve the validity of Chain-CAT, we balanced the content and reduced the exposure of individual items. In the constrained content balancing method, items are selected so that the proportion of concept subgroups in an administered test is similar to that of the original test. Let us assume the case of , for simplicity. The FCI has five subgroups (), as discussed above. Thus, =1 item from each subgroup is administered in each test. If we administer the test (=9 in our simulation) times throughout the duration of the course, the total number of items to administer from each subgroup, , is thus given by times , which equals 9.
@@
For the cases where the test length is longer than 5 items, the calculation of is more complicated because the frequency of administering items is no longer the same for each subgroup. We determined by manually creating a table to lay out the order in which subgroups would be chosen for each subsequent item. We then counted the number of items administered from each subgroup and derived that , and are the minimum number of administrations of a given item without resulting in subgroups becoming unrepresented in tests administered later in the semester. We confirmed these thresholds using numerical simulations. Note that for , means that there is no limit because it equals the number of tests, . That is, for , a given FCI item can appear on each of the nine tests.
@@ II.5.4 Calculation of the accuracy and precision
We represented the accuracy and precision by the bias and variance, respectively. Then, we summarized these measures in terms of the mean-square error (MSE). The MSE is defined by the following equation [47], which equals the sum of the variance and the squared bias,
where is the expected value of . In the equation, is the true proficiency given for each time point and is the estimated proficiency level. The MSE was calculated as the average of the squared difference of and using the above 10 000 response data for each time point. The bias is defined by,
@@
We calculated the accuracy and precision of the Chain-CAT and compared them to those obtained when the FCI is conducted in the ordinary pre-post method. The latter was calculated by having all 30 FCI items administered both in the pretest and post-test and estimating student proficiency using item response theory in the same model as above (2PL model) without collateral information. Specifically, we calculated the RMSE of the pre-post method and used this as a reference value with which to examine the efficiency of the Chain-CAT.
When comparing the accuracy and precision of the Chain-CAT with the pre-post method, we use two different types of averages for the RMSE. The first one is the time-averaged RMSE, which is defined by
where is the initial time point, is the final time point, and is the number of the time points (number of tests, as above). The second type of average RMSE we calculated is the overall average of the RMSE, which is calculated by taking the average of the time-averaged RMSE for all progression models that we assumed in the simulations:

### v2

@@ Abstract
Although conceptual assessment tests are commonly administered at the beginning and end of a semester, this pre-post approach has inherent limitations. Specifically, education researchers and instructors have limited ability to observe the progression of student conceptual understanding throughout the course. Furthermore, instructors are limited in the usefulness of the feedback they can give to the students involved. To address these challenges, we propose an alternative approach that leverages computerized adaptive testing (CAT) and increasing the frequency of CAT-based assessments during the course, while reducing the test length per administration, thus keeping or decreasing the total number of test items administered throughout the course. The feasibility of this idea depends on how far the test length per administration can be reduced without compromising the test accuracy and precision. Specifically, the overall test length is desired to be shorter than when the full assessment is administered as a pretest and subsequent post-test. To achieve this goal, we developed a CAT algorithm that we call Chain-CAT. This algorithm sequentially links the results of each CAT administration using collateral information. We developed the Chain-CAT algorithm using the items of the Force Concept Inventory (FCI) and analyzed the efficiency by numerical simulations. We found that collateral information significantly improved the test efficiency, and the overall test length could be shorter than the pre-post method. Without constraints for item balancing and exposure control, simulation results indicated that the efficiency of Chain-CAT is comparable to that of the pre-post method even if the length of each CAT administration is only 5 items and the CAT is administered 9 times throughout the semester. However, when these constraints are imposed, we found that the efficiency of the Chain-CAT with the test length of 5 items is lower than that of the pre-post method. Based on this result and analysis of which items were administered, we suggest expanding the FCI item bank by creating additional items or incorporating items from other research-based assessments that have high discrimination parameters and exhibit sufficient variability in their difficulty parameters.
@@ I Introduction
The pre-post method is a standard method to analyze the effectiveness of instruction, but there are limitations to the approach. First, the pre- and post-test results provide only snapshots of students’ understanding at the beginning and end of the course, limiting the ability to observe the progression of conceptual understanding throughout the duration of the course. Second, students may see little benefit in their taking the assessment when the focus of the assessment is to reflect on the year’s instruction and improve instructional practices for future students, with no feedback to the students who actually take the assessment. It is therefore not surprising that post-test response rates are generally lower than pretest response rates (especially when students are asked to take the assessment outside of class), unless extra credit is offered to maintain student engagement and motivation [11, 12, 13].
To address these issues associated with the pre-post method, we propose an alternative approach which increases the frequency of the assessments during the course while decreasing the test length per administration, thus maintaining or decreasing the total number of test items during the course. This idea is inspired by the microgenetic method [14, 15, 16, 17, 18, 19, 20, 21], which involves frequent quantitative and/or qualitative measurements. The microgenetic method involves 1) sampling data at a frequency that is assumed to be high compared to the rate of change of the phenomenon of interest and 2) collecting data for the entire period of change [20], allowing teachers to monitor the progression of students’ understanding. Although these methods are generally implemented with small student cohorts, Sayre and Heckler [16, 17] used the microgenetic approach in frequently administering simple multiple choice conceptual questions to a large number of students.
While the pre-post method aligns with diagnostic and summative assessment, our idea of frequent administration of short assessments can be used as a form of formative assessment. In terms of research instruments, the method of survey administration chosen should depend upon the research questions being investigated. In terms of enhancing instruction, our method opens the door to providing automated feedback according to each student’s set of responses. We expect that this will increase the usefulness of the survey for students and reduce their sense of burden in completing it.
The feasibility of the approach we are proposing depends on how far the test length per administration can be reduced. A reference value of the total test length during a course is 60 items, as this is the total test length if the full FCI is administered both as the pre- and post-test. If the course involves 10 administrations of the shortened FCI, it is then desired that the test length per administration would be less than 6 items.
@@
We utilize a CAT algorithm that takes advantage of collateral information [34, 22]. Collateral information refers to relevant empirical data about respondents, such as age, grade level, or previous test scores. In our study, we specifically used only the proficiency estimate from each respondent’s previous test—that is, the most recent proficiency estimate carried over, as is commonly done (e.g., in Ref. [22]). This information can be used to select the first item in CAT, accelerating the convergence of the proficiency estimation, and hence improving test efficiency [22]. Specifically, as we describe below, we use the proficiency estimate of each respondent in a test for selecting items and estimating respondent proficiency level in the next test for the respondent. In contrast to one-time testing, these subsequent tests constitute multiple measurements over time (e.g., with intervals of about a week between them). Since this CAT algorithm sequentially links the test result of each administration, we call this algorithm Chain-CAT. In CAT, the proficiency estimate is updated after each item and used to select the next item at a given point in time using a specified algorithm. In addition to this, Chain-CAT uses collateral information, employing the same algorithm with the proficiency estimate from the last question of the previous test (e.g., one week before) to initiate item selection in the current test. Although the algorithm is the same, the term “collateral information” is introduced to emphasize the time gap between tests, during which learning can occur.
@@
Several researchers have studied the development of a CAT-based formative assessment system. Veldkamp et al. [35] discussed the viability of a computerized adaptive learning system which responds to students based upon their test results. Choi and McClenen [36] developed and analyzed the validity and efficiency of an adaptive formative assessment system based on CAT with the time series proficiency estimation utilizing Bayesian networks. Yang et al. [37] developed a formative assessment system based on CAT and learning memory cycle. While there have been previous studies such as these, our goal is to construct an algorithm from our design of chaining collateral information as described above and to introduce its practical application in the context of physics education research.
The objectives of Heckler and Sayre [17] and our Chain-CAT are partially aligned, particularly in terms of tracking student understanding. However, the methods employed differ significantly. Heckler and Sayre [17] repeatedly administered individual questions to randomly selected groups of students and analyzed changes in the correct response rate of each question. In contrast, our approach employs a CAT-based system in which the test consists of multiple questions. The selection of those questions varies depending on student’s responses and the analysis is based on the proficiency estimates derived from those responses.
The objective of this study is to examine the feasibility of the Chain-CAT version of the FCI (Chain-CAT FCI). Specifically, our research questions are: (i) How much does collateral information improve the test efficiency when it is used to sequentially link the results of each CAT administration? (ii) Is it possible to shorten the total test length of Chain-CAT FCI in a course to be less than that of the full FCI conducted as the pre- and post-test (60 items) without compromising test accuracy and precision? Under what conditions is this possible? (iii) Do the 30 FCI items comprise a sufficient item bank for the effective use of Chain-CAT? If not, what improvement is necessary? The focus of this paper is on the analysis of the feasibility of the Chain-CAT FCI using numerical simulations. The development of feedback methods to deliver formative assessment will be the subject of future work.
@@ II.1 Item response model and item bank
We estimated the item parameters using the FCI survey data obtained in our previous study [43, 44]. In that study, we administered the full-length paper-and-pencil FCI in class to 2882 Japanese university students from April 2015 to April 2018. The examinees were students at the beginning of introductory physics courses at one public and four private universities, all of which are considered mid-ranked institutions in Japan. From this dataset, we removed aberrant responses to be left with 2712 valid responses. The result of the item parameter estimation is shown in Table 3 in the appendix. Although the standard error of the difficulty parameter for question 29 is relatively large, this item was rarely administered in the simulation due to its low discrimination. Hence, we expect that its inclusion has no significant effect on our results.
The assumptions of unidimensionality, overall local independence, and goodness of fit were examined in our previous studies [43, 44], supporting the validity of using the 2PL model in our analysis. Specifically, we assessed the unidimensionality of the FCI using principal component analysis with a tetrachoric correlation matrix [42], and evaluated the assumption of local independence using Yen’s statistic [45]. The results indicated that our FCI dataset has sufficient unidimensionality and local independence at the overall test level, although higher unidimensionality and local independence are desirable (see Sec. IV.2).
We evaluated the goodness of fit of 1PL, 2PL, and 3PL IRT models to the response data with the standardized root mean square residual (SRMSR) [46]. A value of SRMSR less than 0.05 is considered to indicate that the model is well-fitted. The SRMSR values obtained were 0.079 for the 1PL model, 0.041 for the 2PL model, and 0.041 for the 3PL model. These results suggest that both the 2PL and 3PL models fit the data well, whereas the 1PL model does not.
To select the most appropriate model, we used the Bayesian Information Criterion (BIC) [47]. BIC increases if the deviance of the model from the data increases and if the number of parameters increases; thus, the model with the lowest BIC is the most preferable. Based on our data set, the BIC was for the 1PL model, for the 2PL model, and for the 3PL model. These results indicate that the BIC values for the 2PL and 3PL models are comparable and both are lower than that of the 1PL model.
Given the possibility that future studies may compare their results with those of the present study, it is important to reduce the cost of data collection. Since the recommended sample size for accurately estimating item parameters is approximately 500 for the 2PL model and around 2000 for the 3PL model [42], we selected the 2PL model for our analysis.
@@ II.4 Improvement of the validity of Chain-CAT
Second, we controlled the item exposure for the Chain-CAT FCI. Generally, if the items in an item bank are very informative at the average ability level, they are selected more frequently than the other items [23]. Although a given item does not appear more than once in a given test, the simple Chain-CAT algorithm can result in that item frequently appearing across the tests. If students repeatedly answer certain items each week, they may memorize the content of the items they solved the previous week and answer them without rethinking, which may not reflect their actual proficiency. Therefore, we implemented an algorithm that limits the number of times an item can be administered in a course. Specifically, we used an algorithm whereby once an item is administered a specific number of times in a course (ranging from 3 to 9 items, depending on the test length), it is removed from the list of items that can be administered (described in more detail below).
@@ II.5 Numerical simulation procedure
We conducted a numerical simulation to analyze the efficiency of the Chain-CAT and to search for the optimal settings. The outline of the procedure is as follows. (1) Assume a progression model of the true-value proficiency as a function of session (the unit is taken to be 1-2 weeks) in a course. (2) Generate response data for a given true value of based on the item response model and calculate the estimates of . (3) Calculate the accuracy and precision by the root-mean-square error (RMSE) from the true and the estimated values of and compare it to a reference value. In what follows, we explain the details of the above procedure.
@@ II.5.1 Progression models of the true proficiency level
To analyze the efficiency of the Chain-CAT as generally as possible, we assumed various progression patterns of the true-value proficiency . Specifically, we assumed the following progression models for : a) a stationary model in which is constant across sessions, b) a linear model in which increases linearly across sessions, and c) a step model in which increases significantly only at a certain session. Although it is uncommon in introductory physics courses for student scores on the FCI to decrease, we confirmed that the simulation results of the case when decreases linearly or decreases significantly only at a certain point are similar to the results of b) or c), respectively. We supposed that most of the student’s progressions are one of these patterns or a combination of them. Moreover, we analyzed the test efficiency of the Chain-CAT in various proficiency ranges for each model. For example, in the stationary model, we analyzed the efficiency in the middle students and the high students . We confirmed that the low range yields similar results to the high range , so we do not show the result below. We fixed the number of tests conducted throughout the simulated course to be 9 times. This would be a reasonable number of tests to be administered in a course consisting of 10-15 weeks.
@@ II.5.2 Generation of the response data
We generated response data based upon a Monte Carlo simulation, which is commonly used in the development of CAT [50]. Specifically, we generated the responses for Chain-CAT using the function simulateRespondents of the catR package. In the simulation, the response to a selected item is generated based on the calibrated item parameters of that item and the value of true proficiency at a particular session. For example, in Eq. (1), suppose that the probability of a correct response for an item is calculated to be 0.75 for a given true proficiency. A random number is generated from a uniform distribution within a range of 0 to 1. If the value is 0.75 or less, the generated response is coded as “correct” for that item. If the value is greater than 0.75, then the generated response is “incorrect.” The EAP method is used to estimate the proficiency of the respondent, an item selection method (e.g., MFI criterion) is then used to choose the next item based upon that estimated proficiency, and the process repeats until reaching the predetermined test length. Specifically, we varied the test length to be 5, 10, and 15. Then, for each length, we calculated accuracy and precision.
Note that when the test is administered nine times with test lengths of , , and , the total number of items becomes 45, 90, and 135, respectively. In contrast, the pre-post method uses a total of 60 items. Therefore, only in the case of is the total number of items fewer than in the pre-post method. Our analysis in this paper includes results for . We think such test lengths may be viable for instructors, even if the total number of questions exceeds 60. Limiting each session to 10 questions helps distribute the workload, potentially reducing the perceived burden. Moreover, even if the total number of questions is greater, having multiple small tests throughout the semester allows instructors to provide feedback as part of formative assessment.
To summarize, provided calibrated item parameters and a given true proficiency level, the simulateRespondents function generates an entire set of correct/incorrect responses and calculates estimated proficiencies for a given test length of each Chain-CAT administration. This is done for each session in a course. In this analysis, we generated 10 000 response data based on the same true theta for a given session.
@@ II.5.3 Algorithm to improve the validity of Chain-CAT
As we mentioned above, to improve the validity of Chain-CAT, we balanced the content and reduced the exposure of individual items. In the constrained content balancing method, items are selected so that the proportion of concept subgroups in an administered test is similar to that of the original test. Let us assume the case of , for simplicity. The FCI has five subgroups (), as discussed above. Thus, =1 item from each subgroup is administered in each test. If we administer the test at (=9 in our simulation) sessions throughout the duration of the course, the total number of items to administer from each subgroup, , is thus given by times , which equals 9.
@@
For the cases where the test length is longer than 5 items, the calculation of is more complicated because the frequency of administering items is no longer the same for each subgroup. We determined by manually creating a table to lay out the order in which subgroups would be chosen for each subsequent item. We then counted the number of items administered from each subgroup and derived that , and are the minimum number of times administrated of a given item without resulting in subgroups becoming unrepresented in tests administered later in the semester. We confirmed these thresholds using numerical simulations. Note that for , means that there is no limit because it equals the number of tests, . That is, for , a given FCI item can appear on each of the nine tests.
@@ II.5.4 Calculation of the accuracy and precision
We represented the accuracy and precision by the bias and variance, respectively. Then, we summarized these measures in terms of the mean-square error (MSE). The MSE is a commonly used statistic for evaluating the accuracy and precision of parameter estimation in simulation studies. It is effective when the true proficiency parameters (theta) are known, particularly in the context of CAT [23]. The MSE is defined by the following equation [51], which equals the sum of the variance and the squared bias,
where is the expected value of . In the equation, is the true proficiency given for each session and is the estimated proficiency level. The MSE was calculated as the average of the squared difference of and using the above 10 000 response data for each session. The bias is defined by,
@@
Our previous study [43, 44] indicated that, within the 30 items of the FCI, longer test lengths led to smaller RMSE. Although generally, the higher the accuracy, the better, our approach prioritizes features such as test time and application to formative assessment, and so we look for a degree of accuracy that is comparable to that of the paper-based pre-post method. Specifically, we calculated the RMSE for the Chain-CAT and compared it to the RMSE obtained using the conventional pre-post method with the FCI. In the latter approach, all 30 FCI items were administered in both the pretest and post-test, and student proficiency was estimated using item response theory based on the same 2PL model described above, without collateral information. The RMSE from the pre-post method served as a reference value for evaluating the efficiency of the Chain-CAT.
When comparing the accuracy and precision of the Chain-CAT with the pre-post method, we use two different types of averages for the RMSE. The first one is the session-averaged RMSE, which is defined by
where is the initial session, is the final session, and is the number of the sessions (number of tests, as above). The second type of average RMSE we calculated is the overall average of the RMSE, which is calculated by taking the average of the session-averaged RMSE for all progression models that we assumed in the simulations: